Rose’s garden is in the shape of a trapezoid. If the height of the trapezoid is 16 m, one base is 20 m, and the area is 224 m2, find the length of the other base.
Area of a trapezoid: A=1/2h(B+b)
Hints:
B and b are 2 different variables. B is no = to b
B and b are both bases of the trapezoid. (A trapezoid has 2 bases)
B is 20 m; you must find b.
You may want to use your answer to question 24 above to solve this problem, BUT IT IS NOT ACTUALLY NECESSARY; it just makes it slightly easier.
Be sure to check your answer. Put your answer back into the formula for the area of a trapezoid. Does the garden area still equal 224 m2?

Answer by rajagopalan(158)

area=(1/2)*16*(base1+base2)
given base1=20
let base2 =x
then
area=(1/2)*16*(20+x)=224
20+x=(224*2)/16
20+x=28
x=28=20=8
so the other base is 8 m long.
verification:
area=1/2)*16*(28)=224
ok..good day..
(Ps * is used as multiplication symbol to avoid confusion with base x)